OpenStudy (anonymous):

Humberto invested $20,000 part at 15% and part at 7%. The total yield was $2680. How much did he invest at each rate?

OpenStudy (anonymous):

x: invested in 15% 20000-x: invested in 7% \[2680=15x/100+7(20000-x)/100\]\[268000=15x+140000-7x\]\[128000=8x\]\[x=16000\] $16000 in 15% rate, $4000 in 7% rate

OpenStudy (anonymous):

Thank you so much! :)